OpenStudy (anonymous):

radius=? circumference=20pi area=?

jimthompson5910 (jim_thompson5910):

C = 2pi*r 20pi = 2pi*r 20pi/2pi = r 10 = r r = 10 So the radius is 10 units ----------------------------------- A = pi*r^2 A = pi*10^2 A = 100pi So the exact area is 100pi square units